OpenStudy (anonymous):

Solve 2 log x = log 64.

OpenStudy (raden):

hint : 2 log x = log x^2

OpenStudy (raden):

now you have : log x^2 = log 64

OpenStudy (raden):

next, if given log a = log b then a = b. log x^2 = log 64, then x^2 = 64 solve for x

OpenStudy (raden):

just take the squareroot of 64. so x = sqrt(64) = ..... ? (in this case, x can not be negative. it must be positive)